# Quasinormal ringing and Unruh-Verlinde temperature of the Frolov Black Hole

### v2

@@ Abstract
In this study, we investigate electromagnetic and Dirac field axial-perturbations of a charged regular black hole arising from quantum gravity effects, commonly referred to as the Frolov black hole, a regular (nonsingular) black hole solution. We derive the master wave equations for massless electromagnetic and Dirac perturbations and solve them using the standard Wentzel-Kramers-Brillouin (WKB) method along with Padé Averaging. From these solutions, we extract the dominant and overtone quasinormal mode (QNM) frequencies along with the associated grey-body factors, highlighting the deviations introduced by quantum gravity corrections compared to the classical case of Reissner–Nordström black hole. Furthermore, we analyze the Unruh-Verlinde temperature of this spacetime, providing quantitative estimates of how quantum gravity effects influence both quasinormal ringing and particle emission in nonsingular black hole models.
@@ I Introduction
The paper has been organized in the following manner: In Section (II), we describe the Frolov Black Hole metric, in Section (III), we work out the axial type perturbation equations for the electromagnetic and Dirac field cases, in Section (IV) we describe the WKB method and compute the QNM frequencies for the given perturbations in Tables (1) to (8). In Section (V) we compute the grey-body factors associated with the Frolov BH and in Section (VI) we work out the Unruh-Verlinde temperature and compare how quantum corrections modify the standard results.
@@ II The Frolov Black Hole Metric
Figure 1: The metric function $f(r)$ as a funtion of $r$ for various values of $\alpha_{0}$ and with a fixed $q=0$ Increasing $\alpha_{0}$ shifts the curve upward and modifies the horizon structure.
Figure 2: The metric function $f(r)$ as a funtion of $r$ for various values of $\alpha_{0}$ and with a fixed $q=0.5$ Increasing $\alpha_{0}$ shifts the curve upward and modifies the horizon structure.
Figure 3: The metric function $f(r)$ as a funtion of $r$ for various values of $\alpha_{0}$ and with a fixed $q=1$ Increasing $\alpha_{0}$ shifts the curve upward and modifies the horizon structure.
@@
III The Axial Perturbation equations
Since we are working with axial perturbations, we present below the field equations for the electromagnetic and Dirac field perturbations in the given curved background.
@@ III The Axial Perturbation equations
where $F_{\rho\sigma}=\partial_{\rho}A_{\sigma}-\partial_{\sigma}A_{\rho}$ and $A_{\mu}$ is the electromagnetic four-potential.
@@
where $r_{*}$ is the tortoise coordinate. The effective potential for electromagnetic perturbations is given by
@@
$\gamma^{0}\sqrt{f}\,\frac{\partial\psi}{\partial t}$ $+\sqrt{f}\,\gamma^{1}\left(\frac{\partial}{\partial r}+\frac{1}{r}+\frac{1}{4f}\frac{df}{dr}\right)\psi$
$+\frac{\gamma^{2}}{r}\left(\frac{\partial}{\partial\theta}+\frac{1}{2}\cot\theta\right)\psi+\frac{\gamma^{3}}{r\sin\theta}\,\frac{\partial\psi}{\partial\phi}=0$ (11)
@@
$\gamma^{0}\sqrt{f}\,\frac{\partial\phi}{\partial t}$ $+\sqrt{f}\,\gamma^{1}\left(\frac{\partial}{\partial r}+\frac{1}{r}\right)\phi$
$+\frac{\gamma^{2}}{r}\left(\frac{\partial}{\partial\theta}+\frac{1}{2}\cot\theta\right)\phi+\frac{\gamma^{3}}{r\sin\theta}\,\frac{\partial\phi}{\partial\phi}=0$ (12)
@@
Introducing the tortoise coordinate
$r_{*}=\int\frac{dr}{f(r)},$ (14)
and the ansatz for the Dirac spinor
@@
Here $Y_{\ell}^{m\pm 1/2}(\theta,\varphi)$ are the usual spin-weighted spherical harmonics.
@@ IV WKB method and QNM frequencies
To compute the quasinormal mode frequencies and analyze the stability of the Frolov black hole under axial perturbations, we use the WKB approximation. This semi-analytic technique is particularly well-suited for potentials with a barrier-like structure, as is the case here, where the Regge–Wheeler (RW) potential governs the evolution of axial perturbations (see Fig. 4, 5, 6, and 7). The QNMs are the complex eigenfrequencies of spacetime perturbations: their real part represents the oscillation frequency, while the imaginary part encodes the damping rate of the perturbation.
The evolution of axial gravitational perturbations in the Frolov BH background is governed by the master equation,
@@
The functional form of $Q(x)$ dictates whether perturbations decay (implying stability) or grow over time, making its accurate modeling crucial for probing the dynamical behavior of the perturbed spacetime.
@@
The Schrödinger-like equation in curved spacetime was first solved using the WKB method by Schutz and Will in 1985 Schutz_1985 . Iyer and Will subsequently extended the formalism to third order in 1987 Iyer_1987 , significantly improving its precision. Later, Konoplya extended the method up to sixth order Konoplya:2019hlu , and more recently, Matyjasek and Opala pushed the approximation to the thirteenth order Matyjasek:2017psv . While higher-order WKB schemes typically increase accuracy, they may also introduce numerical instabilities. In particular, results obtained at very high orders, such as the thirteenth, sometimes deviate significantly from those at lower orders due to error amplification and lack of convergence Hatsuda:2019eoj .
@@
In this work, we focus primarily on the sixth and eighth order WKB approximations, which strike a good balance between accuracy and computational simplicity.
The sixth-order WKB approximation is expressed as
$i\frac{\omega^{2}-V_{0}}{\sqrt{-2V_{0}^{\prime\prime}}}=\sum_{i=2}^{6}\Lambda_{i},$ (26)
where $n=0,1,2,\dots$ represents the overtone number.
Higher-order WKB formulations, such as the seventh and eighth orders, incorporate increasingly large numbers of terms—616 for $\Lambda_{7}$ and 1,215 for $\Lambda_{8}$—making direct computation cumbersome. Nevertheless, they have been shown to yield improved agreement with full numerical solutions in many cases.
Boundary conditions: The computation of QNM frequencies reduces to solving an eigenvalue problem for the perturbation equation, subject to well-defined boundary conditions. These conditions are chosen to represent a physically meaningful response of the spacetime to a transient disturbance. Therefore, we impose
$\psi(r_{*})\sim e^{\pm i\omega r_{*}},\qquad r_{*}\to\pm\infty,$ (27)
where $r_{*}$ is the tortoise coordinate. The choice $e^{+i\omega r_{*}}$ at $r_{*}\to+\infty$ corresponds to a purely outgoing wave at spatial infinity, while $e^{-i\omega r_{*}}$ at $r_{*}\to$-$\infty$ represents a purely ingoing wave at the event horizon. Physically, these conditions encode the idea that no radiation can emerge from behind the horizon and that all perturbations detected far away are purely outgoing. They describe the characteristic “ringdown” stage of a black hole’s response, i.e, the phase that follows after any external source of perturbation has ceased Kokkotas:1999bd .
@@
As we can see, Table (1) lists the fundamental-mode QNMs for the massless electromagnetic perturbation ($l=2$) computed with 6th- and 8th-order WKB with Padé averaging. Both WKB orders produce very similar values for $Re(\omega)$ and $Im(\omega)$, indicating good convergence of the method for these parameters. As the charge $q$ or the BH parameter $\alpha_{0}$ is increased, the real part of the frequency shifts upward and the modes oscillate faster while the magnitude of the imaginary part decreases, i.e., modes begin to damp more slowly, and thus the damping times lenghten. The table therefore demonstrates that introducing charge and parameter stiffen the effective potential and prolong the decay time of the fundamental electromagnetic mode.
@@
In Table (3) we compute the fundamental electromagnetic QNMs for $l=3$. Compared with the $l=2$ fundamental (Table I), higher multipole index results in larger $Re(\omega)$ and a larger magnitude of $Im(\omega)$, consistent with a steeper and narrower effective potential at larger $l$. The dependence on $q$ and $\alpha_{0}$ follows the same pattern seen before: raising either parameter increases the oscillation frequency and reduces the damping magnitude. Agreement between the 6th- and 8th-order Padé-averaged WKB remains close, supporting numerical stability for these higher multipoles in the parameter ranges shown. The first overtone for $l=3$ in Table (4) again shows the canonical overtone features, it is more strongly damped and typically has a somewhat different real and imaginary parts compared to $n=0$. Parameter trends with respect to $q$ and $\alpha_{0}$ remain consistent with the the analysis of the previous tables.
In Table (5) we list fundamental modes for massless Dirac perturbations with $|k|=2$. Qualitatively the Dirac spectrum mirrors the electromagnetic case: increasing $q$ or $\alpha_{0}$ pushes $Re(\omega)$ upward and makes $Im(\omega)$ less negative, i.e. oscillations become faster and longer-lived as the parameters increase. The 6th- and 8th-order WKB–Padé results are in close agreement for the fundamental Dirac mode, indicating that the chosen WKB orders are sufficiently accurate for these modes and parameter ranges, and similarly, we have Table (6) where the first overtone of the Dirac field at $|k|=2$ shows that overtone pattern, i.e, the real part of the frequencies is slighly lower than $n=0$ case. The parameter dependence is consistent with the other tables; both $q$ and $\alpha_{0}$ reduce damping magnitude and increase oscillation frequency. Practically, this means Dirac overtones also become longer lived and somewhat higher in frequency with increased charge or the BH parameter.
@@
The first overtone for $|k|=3$ in Table (8) exhibits the standard behavior relative to the fundamental: increased damping and a characteristic shift of the real part consistent with overtone structure. Here it is observed that increasing $q$ or $\alpha_{0}$ slightly lowers the oscillation frequency and reduces the damping magnitude. Differences between 6th- and 8th-order results are slightly larger here than for fundamentals, but do not change the qualitative conclusions drawn from the table.
@@ V Grey-body Factors
Grey-body factors quantify the fraction of the initial Hawking radiation that successfully transmits through the effective potential barrier surrounding the black hole, rather than being reflected back toward the event horizon. To compute this quantity, we apply Hawking’s semiclassical formula, now modified with a grey-body factor, to estimate the radiation flux that reaches a distant observer. This approach is particularly relevant when considering the late stages of black hole evaporation, even when using the metric function (Eq. (2)).
As discussed in the literature, the contribution of gravitons to the total radiation flux is negligible — for instance, in the Schwarzschild case, gravitons account for less than $2\%$ of the emitted radiation Page:1976df . Consequently, grey-body factors for test fields are sufficient to characterize the radiation spectrum. Moreover, these factors can be more influential than the Hawking temperature itself in determining the intensity of the observed flux Konoplya:2019ppy .
@@
Because the effective potential has a single barrier-like peak and falls off monotonically toward both infinities, the WKB approximation Konoplya:2003ii can be reliably applied to compute $R$ and $T$. Since $\omega^{2}$ is real, the first-order WKB approximation yields real values for $R$ and $T$, satisfying
@@
$|A_{\ell}|^{2}=1-|R_{\ell}|^{2}=|T_{\ell}|^{2},$ (30)
where $A_{\ell}$ represents the grey-body factor.
@@
$R=\left(1+e^{-2i\pi K}\right)^{-1/2},$ (31)
where $K$ is determined from the equation
$K-i\frac{\omega^{2}-V_{\text{max}}}{\sqrt{-2V^{\prime\prime}_{\text{max}}}}-\sum_{i=2}^{6}\Lambda_{i}(K)=0.$ (32)
Here, $V_{\text{max}}$ and $V^{\prime\prime}_{\text{max}}$ are the value and second derivative of the effective potential at its peak, while $\Lambda_{i}(K)$ denote the higher-order WKB correction terms. It is worth noting that the WKB expansion is asymptotic rather than convergent, and there usually exists an optimal order at which the accuracy is maximal. The appropriate order depends sensitively on the shape of the effective potential. Here we used the 6th6{th} order approximation and $3^{rd}$ order approximation for calculating grey-body factors.

### v3

@@ Abstract
In this study, we investigate electromagnetic and Dirac test field perturbations of a charged regular black hole arising from quantum gravity effects, commonly referred to as the Frolov black hole, a regular (nonsingular) black hole solution. We derive the master wave equations for massless electromagnetic and Dirac perturbations and solve them using the standard Wentzel-Kramers-Brillouin (WKB) method along with Padé Averaging. From these solutions, we extract the dominant and overtone quasinormal mode (QNM) frequencies along with the associated grey-body factors, highlighting the deviations introduced by quantum gravity corrections compared to the classical case of Reissner–Nordström black hole. Furthermore, we analyze the Unruh-Verlinde temperature of this spacetime, providing quantitative estimates of how quantum gravity effects influence both quasinormal ringing and particle emission in nonsingular black hole models.
@@ I Introduction
The paper has been organized in the following manner: In Section (II), we describe the Frolov Black Hole metric, in Section (III), we work out the perturbation equations for the electromagnetic and Dirac field cases, in Section (IV) we describe the WKB method and compute the QNM frequencies for the given perturbations in Tables (1) to (8). In Section (V) we compute the grey-body factors associated with the Frolov BH and in Section (VI) we work out the Unruh-Verlinde temperature and compare how quantum corrections modify the standard results.
@@ II The Frolov Black Hole Metric
Figure 1: The metric function $f(r)$ as a function of $r$ for various values of $\alpha_{0}$ and with a fixed $q=0$ Increasing $\alpha_{0}$ shifts the curve upward and modifies the horizon structure.
Figure 2: The metric function $f(r)$ as a function of $r$ for various values of $\alpha_{0}$ and with a fixed $q=0.5$ Increasing $\alpha_{0}$ shifts the curve upward and modifies the horizon structure.
Figure 3: The metric function $f(r)$ as a function of $r$ for various values of $\alpha_{0}$ and with a fixed $q=1$ Increasing $\alpha_{0}$ shifts the curve upward and modifies the horizon structure.
@@
III Perturbation equations
The parameter $q$ appearing in the metric need not be interpreted uniquely as an electric charge; it may instead be associated with an effective or topological charge arising from non-electromagnetic sources, which justifies treating the electromagnetic field considered here as a test field. Therefore, we present below the field equations for the test electromagnetic and Dirac field perturbations in the given curved background.
@@ III Perturbation equations
where $F_{\rho\sigma}=\partial_{\rho}A_{\sigma}-\partial_{\sigma}A_{\rho}$ and $A_{\mu}$ is the electromagnetic four-potential. It is worth noting that electromagnetic perturbations are treated in the test-field approximation, with backreaction effects and coupled gravito-electromagnetic perturbations neglected.
@@
where $r_{*}$ is the tortoise coordinate, which is defined as:
$r_{*}=\int\frac{dr}{f(r)},$ (7)
The effective potential for electromagnetic perturbations is given by
@@
$\frac{\gamma^{0}}{\sqrt{f}}\,\frac{\partial\psi}{\partial t}$ $+\sqrt{f}\,\gamma^{1}\left(\frac{\partial}{\partial r}+\frac{1}{r}+\frac{1}{4f}\frac{df}{dr}\right)\psi$
$+\frac{\gamma^{2}}{r}\left(\frac{\partial}{\partial\theta}+\frac{1}{2}\cot\theta\right)\psi+\frac{\gamma^{3}}{r\sin\theta}\,\frac{\partial\psi}{\partial\varphi}=0$ (12)
@@
$\frac{\gamma^{0}}{\sqrt{f}}\,\frac{\partial\phi}{\partial t}$ $+\sqrt{f}\,\gamma^{1}\left(\frac{\partial}{\partial r}+\frac{1}{r}\right)\phi$
$+\frac{\gamma^{2}}{r}\left(\frac{\partial}{\partial\theta}+\frac{1}{2}\cot\theta\right)\phi+\frac{\gamma^{3}}{r\sin\theta}\,\frac{\partial\phi}{\partial\varphi}=0$ (13)
@@
Invoking the tortoise coordinate from eqn (7) and the ansatz for the Dirac spinor
@@
Here $Y_{\ell}^{m\pm 1/2}(\theta,\varphi)$ are the usual spin-weighted spherical harmonics. We proceed with the separation of variables by using a property of the spinor spherical harmonics. The angular operator appearing in eqn.(13) acts on $\chi^{\pm}_{jm}$ as an eigenvalue operator. In particular, the spinor harmonics satisfy the following relation, which allows the angular dependence of the Dirac equation to be separated:
@@ IV WKB method and QNM frequencies
The computation of QNM frequencies reduces to solving an eigenvalue problem for the perturbation equation, subject to well-defined boundary conditions. These conditions are chosen to represent a physical response of the spacetime to a transient disturbance. We impose
$\psi(r_{*})\sim e^{\pm i\omega r_{*}},\quad r_{*}\to\pm\infty,$ (24)
where $r_{*}$ is the tortoise coordinate. The choice $e^{+i\omega r_{*}}$ at $r_{*}\to+\infty$ corresponds to a purely outgoing wave at spatial infinity, while $e^{-i\omega r_{*}}$ at $r_{*}\to$-$\infty$ represents a purely ingoing wave at the event horizon. Physically, these conditions encode the idea that no radiation can emerge from behind the horizon and that all perturbations detected far away are purely outgoing. They describe the characteristic “ringdown” stage of a black hole’s response, which is the phase that follows after any external source of perturbation has ceased Kokkotas:1999bd . These boundary conditions select a discrete set of complex frequencies $\omega=\omega_{R}+i\omega_{I}$, where $\omega_{R}$ represents the oscillation frequency and $\omega_{I}<0$ encodes the damping rate.
The evolution of the test field perturbations in the Frolov BH background is governed by the master equation,
@@
The functional form of $Q(x)$ influences the structure of the effective potential in the master equation; the actual decay or growth of perturbations is determined only after specifying the appropriate initial data and boundary conditions which are discussed above.
To compute the QNM frequencies and analyze the stability of the Frolov black hole under given test field perturbations, we use the WKB approximation. This semi-analytic technique is particularly well-suited for potentials with a barrier-like structure, as is the case here, where the Regge–Wheeler (RW) potential governs the evolution of perturbations (see Fig. 4, 5, 6, and 7).
@@
The Schrödinger-like equation in curved spacetime was first solved using the WKB method by Schutz and Will in 1985 Schutz_1985 . Iyer and Will subsequently extended the formalism to third order in 1987 Iyer_1987 , significantly improving its precision. Later, Konoplya extended the method up to sixth order Konoplya:2003ii . While higher-order WKB schemes generally improve accuracy, the WKB expansion itself is asymptotic and may fail to converge at higher orders. To mitigate this limitation, one typically employs Padé approximants, following Matyjasek and Opala  Matyjasek:2017psv , which enhance the convergence properties of the WKB series. It is worth noting that the results obtained at very high orders, such as the thirteenth, sometimes deviate significantly from those at lower orders due to error amplification and lack of convergence Hatsuda:2019eoj .
@@
In this work, we focus primarily on the sixth and eighth-order WKB approximations, which strikes a good balance between accuracy and computational simplicity. The general WKB expression for the frequencies can be expressed as an expansion around the eikonal limit, as follows
$\omega^{2}=$ $V_{0}+A_{2}(K^{2})+A_{4}(K^{2})+A_{6}(K^{2})+\cdots$ (27)
$\hskip-22.76219pt-iK\sqrt{-2V_{0}^{\prime\prime}}\Bigl(1+A_{3}(K^{2})+A_{5}(K^{2})+A_{7}(K^{2})+\cdots\Bigr)$
where the matching conditions for the QNM imply a quantization condition,
$K=n+\frac{1}{2},\qquad n=0,1,2,\ldots$ (28)
with $n$ being the overtone number. Here, $V_{0}$ denotes the value of the effective potential at its maximum, $V_{0}^{\prime\prime}$ represents the second derivative of the potential at this point with respect to the tortoise coordinate, and $A_{i}$ for $i=2,3,4,\ldots$ signify the $i$th-order WKB correction terms.
To improve the accuracy of higher-order WKB calculations, we follow the Padé averaging procedure presented in Ref. Konoplya:2019hlu . Within this approach, one introduces a WKB polynomial $P_{k}(\varepsilon)$ by inserting a formal order-counting parameter $\varepsilon$ into the right-hand side of the WKB expansion, such that
$P_{k}(\varepsilon)=$ $V_{0}+A_{2}(K^{2})\,\varepsilon^{2}+A_{4}(K^{2})\,\varepsilon^{4}+A_{6}(K^{2})\,\varepsilon^{6}+\cdots$ (29)
$\hskip-19.91692pt-iK\sqrt{-2V_{0}^{\prime\prime}}\left(\varepsilon+A_{3}(K^{2})\,\varepsilon^{3}+A_{5}(K^{2})\,\varepsilon^{5}+\cdots\right),$
where the polynomial order $k$ coincides with the order of the WKB approximation, and a formal bookkeeping parameter $\varepsilon$ is introduced to organize the WKB expansion as an asymptotic series and to keep track of successive orders in the WKB approximation. The squared quasinormal frequency is obtained by setting $\varepsilon=1$,
$\omega^{2}=P_{k}(1).$ (30)
We consider a family of the Padé approximants $P^{\tilde{n}}_{\tilde{m}}(\varepsilon)$ for the polynomial $P_{k}(\varepsilon)$ near $\varepsilon=0$ with $\tilde{n}+\tilde{m}=k$,
$P^{\tilde{n}}_{\tilde{m}}(\varepsilon)=\frac{Q_{0}+Q_{1}\varepsilon+\cdots+Q_{\tilde{n}}\varepsilon^{\tilde{n}}}{R_{0}+R_{1}\varepsilon+\cdots+R_{\tilde{m}}\varepsilon^{\tilde{m}}}$, (31)
so that $P^{\tilde{n}}_{\tilde{m}}(\varepsilon)-P_{k}(\varepsilon)=\mathcal{O}(\varepsilon^{k+1})$. The squared frequency is approximated by evaluating the rational function at $\varepsilon=1$,
$\omega^{2}\approx P^{\tilde{n}}_{\tilde{m}}(1).$
[…]
To form a reasonable estimate for order $k$ we use the (ad-hoc) averaging recipe:
  1. 1.
    Compute the set of Padé results $\omega_{\tilde{n}/\tilde{m}}$ with $\tilde{n}+\tilde{m}=k$.
  2. 2.
    Define the central value $\omega_{k}^{(c)}$ (roughly $\tilde{n}\approx\tilde{m}$) and the mean of central values $\omega_{k}^{(m)}$.
  3. 3.
    Identify the $r=\lfloor(k+1)/3\rfloor$ pairs of closest values and compute the averages $\omega_{k}^{(1)},\omega_{k}^{(2)},\ldots$.
  4. 4.
    Adopt as the order-$k$ estimate $\omega_{k}$ the mean of those $r$ closest values and estimate the error by the sample standard deviation $S_{k}$ of the same set.
This procedure typically improves the accuracy compared to the non-Padé WKB result and provides an internal error estimate via. For details, explicit formulas, examples and caveats (including cases where some Padé approximants produce outliers and when extra turning points spoil the WKB applicability) see Konoplya:2019hlu .
@@
As we can see, Table (1) lists the fundamental-mode QNMs for the massless electromagnetic perturbation ($l=2$) computed with 6th- and 8th-order WKB with Padé averaging. Both WKB orders produce very similar values for $Re(\omega)$ and $Im(\omega)$, indicating excellent agreement between the two orders of the method for these parameters. As the charge $q$ or the BH parameter $\alpha_{0}$ is increased, the real part of the frequency shifts upward and the modes oscillate faster while the magnitude of the imaginary part decreases, i.e., modes begin to damp more slowly, and thus the damping times lengthen. The table therefore demonstrates that introducing charge and parameter stiffen the effective potential and prolong the decay time of the fundamental electromagnetic mode.
@@
‘In Table (3) we compute the fundamental electromagnetic QNMs for $l=3$. Compared with the $l=2$ fundamental (Table 1), higher multipole index results in larger $Re(\omega)$ and a larger magnitude of $Im(\omega)$, consistent with a steeper and narrower effective potential at larger $l$. The dependence on $q$ and $\alpha_{0}$ follows the same pattern seen before: raising either parameter increases the oscillation frequency and reduces the damping rate. Agreement between the 6th- and 8th-order Padé-averaged WKB remains close, supporting numerical stability for these higher multipoles in the parameter ranges shown. The first overtone for $l=3$ in Table (4) again shows the overtone features, it is more strongly damped and typically has a somewhat different real and imaginary parts compared to $n=0$. Parameter trends with respect to $q$ and $\alpha_{0}$ remain consistent with the the analysis of the previous tables.
In Table (5) we list fundamental modes for massless Dirac perturbations with $|k|=2$. Qualitatively the Dirac spectrum mirrors the electromagnetic case: increasing $q$ or $\alpha_{0}$ pushes $Re(\omega)$ upward and decreases the damping rate $Im(\omega)$ in magnitude, i.e. oscillations become faster and longer-lived as the parameters increase. The 6th- and 8th-order WKB–Padé results are in close agreement for the fundamental Dirac mode, indicating that the chosen WKB orders are sufficiently accurate for these modes and parameter ranges, and similarly, we have Table (6) where the first overtone of the Dirac field at $|k|=2$ shows that overtone pattern, i.e, the real part of the frequencies is slightly lower than $n=0$ case. The parameter dependence is consistent with the other tables; both $q$ and $\alpha_{0}$ reduce damping magnitude and increase oscillation frequency. Practically, this means Dirac overtones also become longer lived and somewhat higher in frequency with increased charge or the BH parameter.
@@
The first overtone for $|k|=3$ in Table (8) exhibits the standard behavior relative to the fundamental: increased damping and a characteristic shift of the real part consistent with overtone structure. Here it is observed that increasing $q$ or $\alpha_{0}$ increases the oscillation frequency and reduces the damping rate in magnitude. Differences between 6th- and 8th-order results are slightly larger here than for fundamentals, but do not change the qualitative conclusions drawn from the table.
@@ V Grey-body Factors
Grey-body factors quantify the fraction of the initial Hawking radiation that successfully transmits through the effective potential barrier surrounding the black hole, rather than being reflected back toward the event horizon. To compute this quantity, we apply Hawking’s semiclassical formula, now modified with a grey-body factor, to estimate the radiation flux that reaches a distant observer.
As discussed in the literature, the contribution of gravitons to the total radiation flux is negligible, for instance, in the Schwarzschild case, gravitons account for less than $2\%$ of the emitted radiation Page:1976df . Consequently, grey-body factors for test fields are sufficient to characterize the radiation spectrum. Moreover, these factors can be more influential than the Hawking temperature itself in determining the intensity of the observed flux Konoplya:2019ppy .
@@
Because the effective potential has a single barrier-like peak and falls off monotonically toward both infinities, the WKB approximation Konoplya:2003ii can be reliably applied to compute $R$ and $T$. Since $\omega^{2}$ is real, where $R$ and $T$ are the complex reflection and transmission amplitudes, respectively. The normalization condition ensuring energy conservation is
@@
$|\Gamma_{\ell}|^{2}=1-|R_{\ell}|^{2}=|T_{\ell}|^{2},$ (34)
where $\Gamma_{\ell}$ represents the grey-body factor.
@@
$|R|=\left(1+e^{-2i\pi K}\right)^{-1/2}$ (35)
where $K$ is numerically determined from equation (27).
Here, we would like to emphasize that in the QNM problem, the boundary conditions of purely ingoing waves at the event horizon and purely outgoing waves at spatial infinity lead to a discrete set of complex frequencies through a quantization condition. In contrast, the computation of grey-body factors corresponds to a scattering problem with real frequencies, where incident, reflected, and transmitted waves are allowed. In this case, the same WKB expansion evaluated near the maximum of the effective potential determines the parameter $K$ in equation (35).
We employ the 6th6{th} order approximation for calculating grey-body factors without using Padé resummation, because it is not straightforward for grey-body factors, unlike the case of QNMs. The grey-body factors shown in Fig. (8) and Fig. (9) represent the transmission probability $\Gamma(\omega)$ for radiation to tunnel through the curvature-induced potential barrier of the Frolov black hole and reach spatial infinity. In both figures, the curves display the characteristic step-like profile of barrier scattering:
$\Gamma(\omega)\approx 0\quad(\omega\ll V_{\max}^{1/2}),\qquad\Gamma(\omega)\to 1\quad(\omega\gg V_{\max}^{1/2}),$
showing that low-frequency waves are strongly reflected, while high-frequency waves penetrate the barrier almost freely.
A clear multipole hierarchy is visible in both figures. As $\ell$ (electromagnetic case) or $|k|$ (Dirac case) increases, the transmission curves shift toward higher frequencies. Higher multipole moment strengthens the centrifugal contribution to the effective potential, raising and widening the barrier and thus suppressing tunneling at a given frequency.
When comparing the two figures, the Dirac field exhibits a transmission curve that is slightly shifted toward lower frequencies relative to the electromagnetic case. In particular, for comparable multipole indices, the Dirac grey-body factors rise from near zero and approach unity at smaller values of $\omega$. This behaviour originates from the spin-dependent structure of the effective potentials: while the electromagnetic potential is directly proportional to $f(r)\ell(\ell+1)/r^{2}$, the Dirac potential contains additional terms involving derivatives of the metric function. As seen from the effective potential profiles discussed earlier, the peak of the Dirac potential is generally lower than that of the electromagnetic potential for similar angular indices as shown in Figs.(4) to (7). A lower barrier height allows fermionic waves to tunnel through the spacetime more easily, so the transition from reflection-dominated to transmission-dominated regimes occurs at smaller frequencies. Nevertheless, both fields preserve the same qualitative multipole hierarchy and approach the universal high-frequency limit $\Gamma\to 1$, confirming that ultraviolet modes are largely insensitive to both spin and the internal regular core structure.